OpenStudy (idealist10):

How to integrate (2u+1)/(2-2u^2) du?

OpenStudy (idealist10):

\[\frac{ 2u+1 }{ 2-2u^2 }du\]

OpenStudy (anonymous):

factor out the 2 first

OpenStudy (ribhu):

numerator = u+1 + u

OpenStudy (ribhu):

denominator = 2(1-u)(1+u)

OpenStudy (idealist10):

So (1/2) times the integral of (2u+1)/(1-u^2) du?

OpenStudy (ribhu):

1/2(1-u) + u/2(1-u^2)

OpenStudy (ribhu):

mow integrate all this separately. for second integral put 1-u^2 =t and this can be solved.

OpenStudy (idealist10):

Will partial fractions work?

OpenStudy (ribhu):

i don't think they are required nut yes they can be applied in order to get the answer.

OpenStudy (ribhu):

**but

OpenStudy (idealist10):

Okay, thanks!

OpenStudy (ribhu):

ok i am happy that i cud help u